OpenStudy (anonymous):

Write the following quadratic function in the vertex form. Then, identify the axis of symmetry and vertex. y = x2 + 7x - 2

OpenStudy (raden):

it can rewrite as y = x^2 + 7x - 2 = (x + 7/2)^2 - 57/4 do u know why ?

OpenStudy (anonymous):

Not a clue.....Hahah.